“Blast it!” said David Wilson, president of Teledex Company. “We’ve just lost the bid on the Koopers job by $3,000. It seems we’re either too high to get the job or too low to make any money on half the jobs we bid.”

Teledex Company manufactures products to customers’ specifications and uses a job-order costing system. The company uses a plantwide predetermined overhead rate based on direct labor cost to apply its manufacturing overhead (assumed to be all fixed) to jobs. The following estimates were made at the beginning of the year:

Department
Fabricating Machining Assembly Total Plant
Manufacturing overhead $ 371,000 $ 424,000 $ 95,400 $ 890,400
Direct labor $ 212,000 $ 106,000 $ 318,000 $ 636,000

Jobs require varying amounts of work in the three departments. The Koopers job, for example, would have required manufacturing costs in the three departments as follows:

Department
Fabricating Machining Assembly Total Plant
Direct materials $ 4,200 $ 500 $ 2,600 $ 7,300
Direct labor $ 5,200 $ 800 $ 7,400 $ 13,400
Manufacturing overhead ? ? ? ?

Required:

1. Using the company's plantwide approach:

a. Compute the plantwide predetermined rate for the current year.

b. Determine the amount of manufacturing overhead cost that would have been applied to the Koopers job.

2. Suppose that instead of using a plantwide predetermined overhead rate, the company had used departmental predetermined overhead rates based on direct labor cost. Under these conditions:

a. Compute the predetermined overhead rate for each department for the current year.

b. Determine the amount of manufacturing overhead cost that would have been applied to the Koopers job.

4. Assume that it is customary in the industry to bid jobs at 150% of total manufacturing cost (direct materials, direct labor, and applied overhead).

a. What was the company’s bid price on the Koopers job using a plantwide predetermined overhead rate?

b. What would the bid price have been if departmental predetermined overhead rates had been used to apply overhead cost?
Answer #1

Ans.

1 (a)

Plant wide overhead rate for current year = Estimated manufacturing overhead/Estimated direct labor cost

Plant wide overhead rate for current year = $ 890,400 / $ 636,000 = 1.4 or 140% of Direct labor cost

1 (b)

Amount of manufacturing overhead cost that would have been applied to the Koopers job = $ 13,400 * 140% = $ 18,760

2 (a)

Overhead rate for each department:

Fabricating = $ 371,000 / $ 212,000 = 1.75 or 175%

Machining = $ 424,000 / $ 106,000 = 4 or 400%

Assembly = $ 95,400 / $ 318,000 = 0.3 or 30%

2 (b)

Amount of manufacturing overhead cost that would have been applied to the Koopers job.

Fabricating = $ 5,200 * 175% = $ 9,100

Machining = $ 800 * 400% = $ 3,200

Assembly = $ 7,400 * 30% = $ 2,220

So,Total Overhead =  $9,100 + $3,200 + $2,220 = $ 14,250

4 (a)

The company's bid price on the Koopers job if a plantwide overhead rate had been used to apply overhead cost

Total Manufacturing cost = Direct Material + Direct Labor + Manufacuring Overhead (Plantwide)

Total Manufacturing cost = $ 7,300 + $ 13,400 + $ 18,760 = $ 39,460

Company's bid price = (Total Manufacturing cost * 150%) = $ 39,460 * 150% = $ 59,150

4 (b)

The bid price if departmental overhead rates had been used to apply overhead cost

Total Manufacturing cost = Direct Material + Direct Labor + Manufacuring Overhead (Departmental)

Total Manufacturing cost = $ 7,300 + $ 13,400 + $ 14,250 = $ 34,950

Company's bid price = (Total Manufacturing cost * 150%) = $ 34,950 * 150% = $ 52,425